Ariel dropped a golf ball from her second story window. The ball starts from rest

and hits the sidewalk 1.5 s later with a velocity of 14.7 m/s. Find the average

acceleration of the golf ball.



1
Expert's answer
2021-09-29T09:55:03-0400

Given:

"t=1.5\:\rm s"

"v_i=0\:\rm m/s"

"v_f=14.7\:\rm m/s"


The average acceleration of the golf ball

"a=\frac{v_f-v_i}{t}=\frac{14.7-0}{1.5}=9.8\:\rm m/s^2"